1. Find reactions at members KJ, KD and CD (method of joints): (30 marks) AVA 3 m. 3mm 20 KN 30 KN 40 KN
Answer #1

H K E E A С х 40 30 20 sechay X-X cutting fag -FED ofes 30 Atsot 20 56-7 EMG co Ayx18–20x15 - 30412 - 40x950 1 As = 56.11n NoEco 56.7 20-30 „xfgco RD=8.38KN Efxco -25-1+ 3x 8:18 + ferro Fo=70.1 KN fty =-75.1KN (compy fig= 8.28 kN (Pausial fp = 7o. I